OpenStudy (anonymous):

Can you explain me how to solve: 7log9 (x + 8) = 7.

OpenStudy (anonymous):

\[\log9(x+8)=1\]

OpenStudy (anonymous):

\[9(x+8)=e ^{1}\]

OpenStudy (anonymous):

\[9x+72=e ^{1}\]

OpenStudy (anonymous):

can you please explain me the second step ?

OpenStudy (ikram002p):

is 9 is the base of the log or is it just num multipl to (x+8) ?

OpenStudy (anonymous):

it is the base

OpenStudy (ikram002p):

so esam take care of the base e wont help

OpenStudy (anonymous):

the 9 is the base? so my ans/ is not correct

OpenStudy (ikram002p):

yess

OpenStudy (anonymous):

so how can it be solved, can you help me please this topic really confuses me

OpenStudy (ikram002p):

use this hint : log x for base b =y b^y=x

OpenStudy (ikram002p):

7log9 (x + 8) = 7. log9 (x+8)=1 using the hint 9^1=x+8

OpenStudy (ikram002p):

so u have x+8=9 x=1 hope ur getting it nw

OpenStudy (anonymous):

thank you !!

OpenStudy (ikram002p):

np , ur wlc ^^